# KVmix: Gradient-Based Layer Importance-Aware Mixed-Precision Quantization for KV Cache

### v1

@@ 3.3 Dynamic Pivotal Context Selection
In the KV Cache, not all Keys and Values are equally important for generating future tokens. Recent tokens provide the most relevant contextual information for the generation of subsequent tokens, and usually have more impact on the upcoming tokens. We define the Keys and Values corresponding to these pivotal recent tokens as the Recent Pivotal Context (RPC). To optimize model performance while minimizing memory usage, we propose a dynamic RPC selection strategy based on the importance analysis provided by KVmix profiler. Specifically, for the $i$-layer, we assign it an RPC selection ratio $r$ based on the $\bar{s}_{k_{i}}$ and $\bar{s}_{v_{i}}$ scores, with higher $\bar{s}_{k_{i}}$ and $\bar{s}_{v_{i}}$ resulting in larger $r$. The number of RPCs is computed by $num\_RPC=\lfloor r\times current\_RPC\rfloor$. $current\_RPC$ is the sum of the number of new Key and Value states at current time step and the number of historical RPCs.
@@
Figure 3: Dynamic adjustment process of KV Cache quantization in the prefill and decoder phases.
The corresponding number of KV pairs is selected as RPCs based on $num\_RPC$. We keep full precision for the RPCs while performing the mixed quantization for the less critical and older KV pairs. This strategy ensures that the number of full-precision RPCs is dynamically reduced in real-time during long context inference, thus avoiding excessive memory pressure caused by preserving a large number of full-precision KV pairs, while maintaining high-quality sequence generation. In addition, since the importance of Keys and Values may differ within the same layer, the RPC selection ratio for Keys and Values varies within that layer. The RPC selection ratio can be adjusted to provide flexibility in the dynamic balance between inference accuracy and memory usage.
3.4 CUDA Implementation
@@
During model inference, the quantization of the KV Cache introduces additional overhead from quantization and dequantization operations. To mitigate this overhead and enhance inference efficiency, we design efficient CUDA kernels for quantization, dequantization, and matrix-vector multiplication. ① Fusion of quantization and concatenation in the decoding phase. In the decoding phase, the Key states and Value states of the current layer need to be concatenated with the historical KV Cache. Directly quantizing the current states before concatenation incurs additional memory access overhead. Therefore, we fuse the quantization process with the concatenation operation into a single CUDA kernel. Specifically, the kernel processes each element in a streaming manner: each element of the Key states and Value states is quantized and immediately appended to the quantized historical KV Cache, which reduces memory access overhead. We also leverage CUDA thread blocks to process data from different tokens in parallel and utilize shared memory to cache intermediate results, further improving data locality. ② Fusion of dequantization and matrix-vector multiplication in attention computation. During attention computation, the quantized KV Cache must be dequantized before performing matrix-vector multiplication. Dequantizing the entire KV Cache prior to multiplication increases memory usage. To mitigate this issue, we fuse dequantization with matrix-vector multiplication. Specifically, each element is dequantized and immediately multiplied and accumulated with its corresponding element, minimizing additional memory overhead. ③ Efficient kernels for multi-bit quantization configurations. To support the various quantization bit-widths of KVmix, we develop efficient CUDA kernels for 1-bit, 2-bit, 3-bit, and 4-bit quantization. Additionally, we implement customized matrix-vector multiplication kernels for each quantization configuration, ensuring compatibility with varying bit-widths.
4 Experimental Results
4.1 Experimental Setup
@@
We evaluated the proposed method using the Llama 2-7B-hf [32], Llama 3-8b-Instruct, Llama 3.1-8B [33], and Mistral-7B-Instruct-v0.3 [34] models. The datasets were selected based on three distinct evaluation schemes: ① Long Context Evaluation: We used the LongBench [35] benchmark to assess performance on long-context tasks. This benchmark encompasses key long-text application scenarios, including single-document question answering (QA), multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion. The maximum sequence length for all tasks was set to 4096. ② Language Modeling: We measured the perplexity of our method on the Wikitext-2 [36] dataset to evaluate its language modeling capabilities. ③ Mathematical Reasoning: We employed the GSM8K [37] dataset to assess the model’s performance on mathematical reasoning tasks.
We used the NVIDIA RTX 4090 GPU (24GB) to evaluate the quantized model’s inference efficiency and the KV cache’s compression rate. Section 4.2 elaborates on the specific configurations for KV quantization bits and the proportion of RPC.
4.2 Profiling Results
@@
We used 3-bit and 2-bit mixed quantization for Key, and 4-bit and 2-bit mixed quantization for Value. The proportion of 4-bit and 3-bit is 20% of the model layers, and the remaining model layers are all quantized with 2-bit, that is, the average quantization bit of Key is 2.19 (exact value is 2.1875), and the average quantization bit of Value is 2.38 (exact value is 2.375). We randomly selected 30 prompts from the LongBench for KV importance analysis. Using the KVmix profiler, we obtained the KV bit configurations and RPC proportions for each layer. Fig. 4 shows the detailed configurations. When Key and Value are quantized to 3 bits and 4 bits, their RPC proportions are set to 20%; for 2-bit quantization, the RPC proportions are set to 10%. The group size for quantization is 32. The following experiments use this configuration unless otherwise specified. Users can flexibly customize the quantization configuration by modifying the proportion of layers with different bits and the proportion of RPC in KVmix profiler according to different model accuracy or memory requirements.
@@
4.3 Performance Evaluation
4.3.1 Long Context
We evaluated the performance of various quantization configurations across 8 distinct datasets from the LongBench benchmark, using the FP16 model as the baseline. The detailed experimental results are presented in the Table 1. The findings indicate that KVmix-k2.19v2.38 achieves an average accuracy loss (average values on 4 different models) of 0.92% compared to the FP16 baseline. In contrast, the average accuracy loss of KVmix-2bit reached 2.14% relative to the baseline. In addition, random selection of high-bit quantization layers (random-k2.19v2.38) leads to an average accuracy loss of 1.78% compared to the baseline. These accuracy losses are significantly higher than KVmix-k2.19v2.38, demonstrating the advantage of KV importance-aware mixed quantization. Appendix C presents additional evaluation results on LongBench. Increasing the proportion of high-bit quantization layers for KV to 30% (KVmix-k2.28v2.56) significantly improves model accuracy compared to KVmix-k2.19v2.38, nearly matching the performance of full 4-bit quantization. Certainly, compared to KVmix-k2.19v2.38, KVmix-k2.28v2.56 increases memory usage.
Table 1: Model accuracy of 4 LLMs on LongBench with different quantization configurations. KVmix-k2.19v2.38 uses the configurations of Section 4.2. KVmix-2bit uses the asymmetric 2-bit (Key per-channel and Value per-token) quantization for all model layers (RPC ratio is set to 10%). random-k2.19v2.38 randomly selects 20% of the model layers to perform asymmetric 3-bit and 4-bit quantization for Key and Value (RPC ratio is set to 20%), and the remaining layers are 2-bit quantization (RPC ratio is set to 10%).
@@ 4.3.1 Long Context
TriviaQA Qasper MF-en QMSum 2WikiMQA Rbench-P TREC PsgRetr-en
Llama-2-7B-hf FP16 78.89 9.55 22.86 21.19 9.94 55.64 66.00 6.64 33.839
@@
4.3.2 GSM8K and Wikitext-2
We evaluated the capabilities of the quantized model in language modeling and mathematical reasoning using the FP16 model as a baseline. The evaluation was conducted using the lm_eval [38] framework, where the quantized model replaced the Hugging Face model. Specifically, we measured the accuracy on the GSM8K dataset and the perplexity on the Wikitext-2 dataset. The experimental results are detailed in Table 2. The results show that 2bit (k-T, v-T) suffers a catastrophic performance loss on GSM8K and Wikitext-2, and the model almost loses its reasoning ability. For 4bit (k-T, v-T), the performance loss of the model on GSM8K and Wikitext-2 also reached 9.17% and 5.28%, respectively.
Table 2: Model accuracy on GSM8K and perplexity on Wikitext-2 using Llama 2-7B-hf. 2bit (k-T, v-T) uses the symmetric 2-bit (Key per-token and Value per-token) quantization for all model layers, and 4bit (k-T, v-T) uses the symmetric 4-bit quantization; their RPC ratio is set to 0.
Methods GSM8K ($\uparrow$) Wikitext-2 ($\downarrow$)
@@ 4.3.2 GSM8K and Wikitext-2
In contrast, on the Wikitext-2, the perplexity score of KVmix-k2.19v2.38 is almost comparable to the baseline, while on the more challenging GSM8K mathematical reasoning task, KVmix-k2.19v2.38 has an accuracy loss of 2.00%, which significantly outperforms the 2bit (k-T, v-T) and the 4bit (k-T, v-T). Moreover, on GSM8K, KVmix-k2.19v2.38 shows a significant accuracy improvement compared to KVmix-2bit and random-k2.19v2.38, which do not leverage the KV importance analysis for more accurate quantization. These results demonstrate the superior performance of KVmix-k2.19v2.38 in complex task reasoning.
4.4 Inference Efficiency and Memory Usage Evaluation
We evaluated the inference throughput and memory usage of KVmix during inference on an NVIDIA RTX 4090 GPU (24GB). To ensure a fair comparison, we applied identical input data across all evaluated methods. The number of input tokens is 688, the maximum number of new tokens is set to 1024, and the model is Llama 2-7B-hf. We compared KVmix against the methods optimized for memory and throughput: KIVI [13] with 2-bit quantization (full precision residual of 64) and Atom [23] with 4-bit quantization. Memory usage results are illustrated in Fig. 6, with a batch size fixed at 4. The reported memory usage represents the dynamic peak memory consumption during inference minus the memory occupied by the model before inference. To fully utilize the GPU memory, we incrementally increased the batch size to explore KVmix’s maximum inference throughput. The throughput results are shown in Fig. 6. The baseline (FP16), Aotm-4bit, and KIVI-2bit-r64 reach out of memory at batch sizes of 4, 18, and 28, respectively, while the KVmix-k2.19v2.38 can reach a maximum batch size of 30 with an inference throughput of 1032 tokens per second.
The results reveal that KVmix-k2.19v2.38 achieves a 4.9$\times$ reduction in inference runtime memory usage and up to a 5.3$\times$ increase in throughput compared to the baseline. This efficient memory compression stems from KVmix’s extremely low bit quantization and dynamic RPC strategy, which progressively reduces the full-precision KV Cache as inference progresses. In contrast, KIVI employs a fixed full-precision residual strategy, unable to dynamically reduce the number of full-precision KVs. Thus, KVmix saves more memory than KIVI’s 2-bit quantization despite using Key-2.19 and Value-2.38 bit configuration. Meanwhile, Atom-4bit quantizes both model weights and activations while utilizing Tensor Cores for optimized kernel execution, achieving a higher throughput at the same batch size but incurring greater model accuracy degradation, for example, it has an average accuracy loss of nearly 10% using the Llama 2-7B-hf model on the Wikitext-2 dataset.
Figure 5: Dynamic peak memory usage of different methods during inference on the Llama 2-7B-hf model with a batch size of 4.
@@ 4.4 Inference Efficiency and Memory Usage Evaluation
4.5 Comparison with SOTA Methods
We compared KVmix against prior SOTA methods for KV Cache, specifically Key per-channel and Value per-token methods, i.e., KIVI [13] and KVQuant [22], since they are known to minimize KV quantization errors. Table 3 presents the model accuracy comparison on the LongBench. The results show that KVmix-k2.19v2.38 outperforms KIVI-2bit-r64, and reduces the average accuracy loss by 1.57%. Although KVQuant-3bit-1% is competitive with KVmix-k2.19v2.38 in terms of accuracy, it achieves at most 4.8$\times$ memory compression and 1.70$\times$ inference acceleration using Llama 2-7B-hf, which is lower than the 4.9$\times$ memory compression and 5.32$\times$ inference acceleration achieved by KVmix-k2.19v2.38. By increasing the quantization bits of KVmix, the advantage of KVmix-k2.28v2.56 over KVQuant in terms of model accuracy is now obvious. Moreover, KVmix-k2.28v2.56 maintains a 4.8$\times$ memory compression and a 5.23$\times$ inference acceleration. This flexibility in balancing accuracy and quantization bits is also one of the important advantages of KVmix.
Table 3: Accuracy comparison of different quantization methods on LongBench using the Llama 2-7B-hf model. KIVI-2bit-r64 uses 2-bit quantization with a full-precision residual of 64. KVQuant-3bit-1% uses 3-bit quantization and 1% outlier handling. KVmix-k2.28v2.56 increases the proportion of high-bit quantization layers in KVmix-k2.19v2.38 to 30%, its detailed configurations are shown in Appendix C, Fig. 8.
FP16 78.89 9.55 22.86 21.19 9.94 55.64 66.00 6.64 33.839
KIVI-2bit-r64 77.08 9.16 22.55 20.12 9.05 56.15 66.00 5.62 33.193
KVQuant-3bit-1% 78.79 10.51 22.61 20.58 9.75 55.62 66.00 5.76 33.703
KVmix-k2.19v2.38 78.78 9.59 22.82 20.49 9.77 56.54 66.00 5.72 33.714
5 Conclusion

### v2

@@ Introduction
In the KV Cache, not all Keys and Values are equally important for generating future tokens. Recent tokens provide the most relevant contextual information for the generation of subsequent tokens and typically have more impact on the tokens to be generated. We define the KVs corresponding to these pivotal recent tokens as the Recent Pivotal Context (RPC). To optimize model performance while minimizing memory usage, we propose a dynamic RPC selection strategy based on the importance analysis provided by KVmix profiler. Specifically, for the $i$-layer, we assign it an RPC selection ratio $r$ based on the $\bar{s}_{k_{i}}$ and $\bar{s}_{v_{i}}$ scores, with higher $\bar{s}_{k_{i}}$ and $\bar{s}_{v_{i}}$ resulting in larger $r$. The number of RPCs is computed by $num\_RPC=\lfloor r\times current\_RPC\rfloor$. current_RPCcurrent\_RPC is the sum of the number of new KV states at the current time step and the number of historical RPCs. The corresponding number of KV pairs is selected as RPCs based on num_RPCnum\_RPC. We keep full precision for RPCs while performing mixed quantization for less critical and older KV pairs, illustrated in Fig. 4. This strategy ensures that the number of full-precision RPCs is dynamically reduced in runtime during long context inference, thus avoiding excessive memory pressure caused by preserving a large number of full-precision KV pairs, while maintaining high-quality sequence generation. Additionally, since the importance of Key and Value may differ within the same layer, the RPC selection ratio for Key and Value varies accordingly within that layer. The RPC selection ratio can be adjusted to balance accuracy and memory: increasing it improves accuracy but requires more memory.
@@
Figure 4: Dynamic adjustment of quantized KV Cache based on RPC during prefill and decoding phases.
CUDA Implementation
@@
During model inference, the quantization of the KV Cache introduces additional overhead due to quantization and dequantization operations. To improve inference efficiency, we design efficient CUDA kernels for quantization, dequantization, and matrix-vector multiplication. ① Fusion of quantization and concatenation. In the decoding phase, the KV states of current layer are concatenated with the historical KV Cache. Quantizing the current states before concatenation causes extra memory access overhead. We fuse quantization and concatenation into a single CUDA kernel, processing each element in a streaming manner. KV elements are quantized and appended directly to the historical KV Cache, thereby reducing memory access. CUDA thread blocks process tokens in parallel, and shared memory caches intermediate results to enhance data locality. ② Fusion of dequantization and matrix-vector multiplication. In attention computation, the quantized KV requires dequantization before matrix-vector multiplication. Dequantizing the full KV beforehand increases memory usage. We fuse dequantization with multiplication, dequantizing each element on-the-fly and immediately multiplying and accumulating it with its corresponding element, minimizing memory overhead. ③ Efficient kernels for multi-bit quantization configurations. To support KVmix’s various quantization bit-widths, we develop CUDA kernels for 1-, 2-, 3-, and 4-bit quantization, along with tailored matrix-vector multiplication kernels for each configuration, ensuring compatibility across bit-widths.
Experimental Results
Experimental Setup
@@
We evaluated the proposed method using Llama 2-7B-hf, Llama 3-8B-Instruct, Llama 3.1-8B (Grattafiori et al. 2024), Mistral-7B-Instruct-v0.3 (Jiang et al. 2023), and Falcon-7B (Almazrouei et al. 2023) models. The datasets were selected based on three distinct evaluation schemes: ① Long Context Evaluation: We used the LongBench (Bai et al. 2024) benchmark to assess performance on long-context tasks. It encompasses multiple key long-text application scenarios. Due to the limited GPU memory, the maximum sequence length was set to 4096. ② Language Modeling: We measured the perplexity on the Wikitext-2 (Merity et al. 2016) dataset to evaluate its language modeling capabilities. ③ Mathematical Reasoning: We employed the GSM8K (Cobbe et al. 2021) dataset to assess the model’s performance on mathematical reasoning tasks. We used the NVIDIA RTX 4090 GPU (24GB) to evaluate the model’s inference efficiency and the KV cache’s compression rate.
Profiling Results
@@
We used 3-bit and 2-bit mixed quantization for Key, and 4-bit and 2-bit mixed quantization for Value. When the KV is quantized to 3 bits or 4 bits, the RPC proportion is set to 20%; for 2-bit quantization, the RPC proportion is set to 10%. Appendix C shows the impact of different RPC proportions on model performance. When the RPC proportion exceeds 20%, its contribution to accuracy improvement is marginal; thus, we selected 20% as the high-bit configuration for KVmix. The group size for quantization is 32. We selected 30 prompts from the LongBench for KV importance analysis. By the KVmix profiler, we can obtain the KV bit configurations and RPC proportions for each layer. When utilizing the KVmix profiler, randomly selecting 20 to 30 prompts is sufficient to yield reliable importance analysis results (Appendix D provides analysis results derived from different prompts by the KVmix profiler); additional prompts do not significantly alter the outcomes. For the models and experimental environment used in this work, this process requires only 10 to 15 minutes, highlighting the efficiency of the KVmix profiler. Users can flexibly customize the quantization configuration by adjusting the proportion of layers with different bit-widths in the KVmix profiler to meet varying accuracy or memory requirements. Fig. 5 illustrates the trends in accuracy, KV memory usage, and throughput as we varied the proportion of model layers quantized to 3 and 4 bits. When the proportion of model layers quantized to 3 and 4 bits is set to 20%, the optimal tradeoff among these three factors is achieved. Under this configuration, the average quantization bit-width for Key is 2.19 (exact value: 2.1875), and for Value is 2.38 (exact value: 2.375). The detailed configuration obtained using the KVmix profiler is shown in Fig. 6. Unless otherwise specified, the configuration is applied to the k-2.19v2.38 quantization in subsequent experiments.
Figure 5: Performance variation of Llama 2-7B with different quantization configurations ("10%" indicates the top 10% important layers are quantized to 4 and 3 bits, and the remaining layers are quantized to 2 bits. The dataset is GSM8K).
@@
Performance Evaluation
Long Context
We evaluated the performance of various quantization configurations across 8 distinct datasets from the LongBench benchmark, using the FP16 model as the baseline. The detailed experimental results are presented in the Table 1. The findings indicate that KVmix-k2.19v2.38 achieves an average accuracy loss (average values on 4 different models) of 1.67% compared to the FP16 baseline. In contrast, the average accuracy loss of KVmix-2bit reached 4.53% relative to the baseline. In addition, random selection of high-bit quantization layers (random-k2.19v2.38) leads to an average accuracy loss of 4.06%. These accuracy losses are significantly higher than KVmix-k2.19v2.38, demonstrating the advantage of KV importance-aware mixed quantization. The average accuracy of KVmix-k2.19v2.38w/oRPC decreases by 3.28% compared with KVmix-k2.19v2.38, proving the effectiveness of using RPC for improving model accuracy. Appendix E presents the evaluation results on additional datasets and increasing the high-bit quantization layer to 30% (KVmix-k2.28v2.56) on LongBench. KVmix-k2.28v2.56 nearly matches full 4-bit quantization performance but uses more memory than KVmix-k2.19v2.38.
@@
TriviaQA
Qasper
MF-en
QMSum
2WikiMQA
Rbench-P
TREC
PsgRetr-en
Llama-2-7B FP16 78.89 9.55 22.86 21.19 9.94 55.64 66.00 6.64 33.839
@@
Falcon-7B FP16 6.94 3.87 7.47 3.96 4.87 12.92 14.00 3.95 7.248
KVmix-2bit 5.96 3.15 6.19 3.28 4.22 11.40 13.50 3.21 6.364
random-k2.19v2.38 6.11 3.10 6.36 3.26 4.24 11.45 13.50 3.26 6.410
KVmix-k2.19v2.38w/oRPC 6.05 3.01 6.54 3.22 4.13 11.41 13.00 3.22 6.323
KVmix-k2.19v2.38 6.64 3.28 7.06 3.52 4.60 12.71 14.00 3.62 6.929
Table 1: Model accuracy of 4 LLMs on LongBench with different quantization configurations. KVmix-k2.19v2.38 uses the configurations of Fig. 6. KVmix-2bit uses the asymmetric 2-bit (Key per-channel and Value per-token) quantization for all model layers (RPC ratio is set to 10%). random-k2.19v2.38 randomly selects 20% of the model layers to perform asymmetric 3-bit and 4-bit quantization for Key and Value (RPC ratio is set to 20%), and the remaining layers are 2-bit quantization (RPC ratio is set to 10%). KVmix-k2.19v2.38w/oRPC is KVmix-k2.19v2.38 without RPC (RPC ratio is set to 0%).
[…]
FP16 78.89 9.55 22.86 21.19 9.94 55.64 66.00 6.64 33.839
KIVI-2bit-r64 77.08 9.16 22.55 20.12 9.05 56.15 66.00 5.62 33.216
QJL-3bit 78.25 9.10 22.60 20.45 9.68 56.09 66.00 5.72 33.486
KVQuant-3bit-1% 78.79 10.51 22.61 20.58 9.75 55.62 66.00 5.76 33.703
KVmix-k2.19v2.38 78.78 9.59 22.82 20.49 9.77 56.54 66.00 5.72 33.714
Table 2: Accuracy comparison of different quantization methods on LongBench using the Llama 2-7B-hf model. KIVI-2bit-r64 uses 2-bit quantization with a full-precision residual of 64. KVQuant-3bit-1% uses 3-bit quantization and 1% outlier handling. QJL-3bit uses 3-bit quantization. KVmix-k2.28v2.56 increases the proportion of high-bit quantization layers in KVmix-k2.19v2.38 to 30%, its detailed configurations are shown in Appendix E.
GSM8K and Wikitext-2
We evaluated the capabilities of the quantized model in language modeling and mathematical reasoning using the FP16 model as the baseline. Atom (Zhao et al. 2024) exhibits very poor performance in long contexts, and thus, we only compare it in this section. The evaluation was conducted using the lm_eval (Gao et al. 2024) framework, where the quantized model replaced the Hugging Face model. Specifically, we measured the accuracy on the GSM8K dataset and the perplexity on the Wikitext-2 dataset. The experimental results are detailed in Table 3. The results show that 2bit (k-T, v-T) suffers a catastrophic performance loss on GSM8K and Wikitext-2, and the model almost loses its reasoning ability. For 4bit (k-T, v-T), the performance loss of the model on GSM8K and Wikitext-2 also reached 9.17% and 5.28%, respectively. In contrast, on the Wikitext-2, the perplexity score of KVmix-k2.19v2.38 is almost comparable to the baseline, while on the more challenging GSM8K mathematical reasoning task, KVmix-k2.19v2.38 has an accuracy loss of 2.00%, which significantly outperforms the 2bit (k-T, v-T) and the 4bit (k-T, v-T). Moreover, on GSM8K, KVmix-k2.19v2.38 shows a significant accuracy improvement compared to KVmix-2bit and random-k2.19v2.38, which do not leverage the KV importance analysis for more accurate quantization. Compared to Atom-4bit and other SOTA methods, KVmix-k2.19v2.38 also has an accuracy advantage. Notably, the Atom-4bit performs 4-bit quantization on both the model weights and activations, which results in greater accuracy loss. These results demonstrate the superior performance of KVmix-k2.19v2.38 in complex task reasoning.
[…]
@@
Table 3: Model accuracy (acc) on GSM8K and perplexity (ppl) on Wikitext-2 using Llama 2-7B-hf. 2bit (k-T, v-T) uses the symmetric 2-bit (Key per-token and Value per-token) quantization for all model layers, and 4bit (k-T, v-T) uses the symmetric 4-bit quantization; their RPC ratio is set to 0.
Inference Efficiency and Memory Usage Evaluation
We evaluated the inference throughput and memory usage of KVmix during inference. To ensure fairness, we applied identical input data across all evaluated methods. The number of input tokens is 688, the maximum number of new tokens is set to 1024, and the model is Llama 2-7B-hf. We compared KVmix against the KIVI-2bit-r64, KVQuant-3bit-1%, QJL-3bit, and Atom-4bit. Memory usage results are illustrated in Fig. 7, with a batch size fixed at 4. The reported memory usage represents the peak memory usage during inference minus the memory occupied by the model before inference. To fully utilize the GPU memory, we incrementally increased the batch size to explore KVmix’s maximum throughput. The throughput results are shown in Fig. 8. The baseline (FP16), Aotm-4bit, and KIVI-2bit-r64 reach out of memory at batch sizes of 4, 18, and 28, respectively, while the KVmix-k2.19v2.38 can reach a maximum batch size of 30 with an inference throughput of 1032 tokens per second.
Figure 7: Dynamic peak memory usage of different methods during inference on the Llama 2-7B-hf model.
@@
[…]
Conclusion